OpenStudy (drummmlers):

Subtract. 5x + 10/x^2 - 5x -14 minus 4/x-7

OpenStudy (triciaal):



OpenStudy (anonymous):

First, find the common denominator

OpenStudy (anonymous):

which is (x-7)(x+2)

OpenStudy (triciaal):

it will be much faster and less complicated to simplify the first expression first 5x + 10 = 5(x + 2) the quadratic factors as (x-7)(x+2) divide by (x+2) 5/(x-7) - 4/(x-7) = 1/(x-7)